OpenStudy (anonymous):

give me the the iupac name and symbol of elemnts of atomic number 123,126,134,148 and 150

OpenStudy (anonymous):

heh anyone there to help me for thisssssssss

sam (.sam.):

http://www.ptable.com/

OpenStudy (anonymous):

not ablee to understnd from the table if u donot mind can u tell me answer here

sam (.sam.):

Wait, 123,126,134,148 and 150 doens't exist

sam (.sam.):

atomic number or atomic mass?

sam (.sam.):

The last atomic number can only be 118, which is Uuo

OpenStudy (anonymous):

itss atomic no. onlyy as gven in qustion

OpenStudy (anonymous):

and if it is mass then wht is the answers

OpenStudy (anonymous):

Though they don't exist , but we learn how to name them . See : \(\quad \begin{array}{|c|c|c|} \hline \textbf{Digit} & \textbf{Name } & \textbf{Abbreviation } \\ \hline 0 & \textbf{nil} & \mathsf{n}\\ \hline 1 & \textbf{un} & \mathsf{u} \\ \hline 2 & \textbf{bi} & \mathsf{b}\\ \hline 3 & \textbf{tri} & \mathsf{t}\\ \hline 4 & \textbf{quad }& \mathsf{q} \\ \hline 5 & \textbf{pent}& \mathsf{p}\\ \hline 6 & \textbf{hex} & \mathsf{h} \\ \hline 7 & \textbf{sept} & \mathsf{s} \\ \hline 8 & \textbf{oct} & \mathsf{o} \\ \hline 9 & \textbf{enn} & \mathsf{e} \\ \hline \end{array} \) Using the above table, you can name the elements with atomic number above 100. For example : Atomic Number 118 --> \(\) \(\begin{equation*} \left. \begin{aligned} 1 \rightarrow \textbf{Un} \\ 1 \rightarrow \textbf{Un} \\ 8 \rightarrow \textbf{Oct} \\ \end{aligned} \right\} \quad\text{118} \implies \mathsf{ Ununoctium } \end{equation*}\)

OpenStudy (anonymous):

You can similarly find for 123, and others :)

OpenStudy (anonymous):

@kapuria , can u do it now?

OpenStudy (anonymous):

nopessssssss

OpenStudy (anonymous):

I will do one for you. 123 : 1 stands for "un" , 2 stands for "bi" and 3 stands for "tri" . So 123 will be : \(\textbf{unbitrium } \) and the symbol will be : ubt